OpenStudy (anonymous):

Write the solution set for equation 7x - 2y = 8 with this replacement set: {(1,-1),(2,3), (0,-4), (4,10)}. A. {(1,-1),(2,3), (0,-4), (4,10)} B. {(1,-1),(2,3), (0,-4)} C. {(2,3), (0,-4), (4,10)}.

OpenStudy (anonymous):

{(1,-1),(2,3), (0,-4), (4,10)}

OpenStudy (anonymous):

7x - 2y = 8 seperate x and y first 7x -8 = 2y then put 1 in for x to see if y will then equal 1 7(1)= 2y +8 7 - 8 =2y (-1 = 2y) -1/2=y...(1,1) not a solution 7(2) -8= 2y 14-8 = 2y (6 = 2y) /2 3=y (2,3) a valid solution set